OpenStudy (anonymous):

express the complex number -3+2i in trigonometric form with theta (0,2π), thx.

OpenStudy (kc_kennylau):

For any complex number z=a+bi, we can express z as r(cosθ+isinθ), where: \(\Large r=\sqrt{a^2+b^2}\) \(\Large \theta=\tan^{-1}\dfrac ba\)

OpenStudy (anonymous):

Thanks, I kind of have forgotten that over a long semester... That was very helpful.

OpenStudy (kc_kennylau):

You could just solve a=rcosθ and b=rsinθ and you'd obtain the same result.